OpenStudy (anonymous):

Given the linear correlation coefficient r and the sample size n, determine the critical values of r and use your finding to state whether or not the given r represents a significant linear correlation. Use a significance level of 0.05. r = 0.843, n = 5

OpenStudy (dumbcow):

using t-test, the test statistic is \[t = \frac{r \sqrt{n-2}}{\sqrt{1-r ^{2}}}\] t = 2.71 at alpha of .05, assuming 2 tailed test critical value is 3.182 conclude it does not represent significant correlation

OpenStudy (dumbcow):

Also critical values for r +- .8783